For the same firm, what is the average total cost?

A firm produces 200 units hourly with $150 in labor costs an hour, $80 in materials costs an hour, and a fixed cost that works out to $20 per hour.

a. $1.15

b. $1.25

c. $250

d. $1
Answer #1

A firm produces 200 units hourly with $150 in labor costs an hour, $80 in materials costs an hour, and a fixed cost that works out to $20 per hour.

Q=200

Total variable cost=$150+$80

Fixed cosr=$20

Total cost=TVC+TFC

=230+20

TC=$250

Average total cost=TC/Q

=250/200

=1.25

Average total cost=$1.25